OpenStudy (anonymous):

The line with slope 5 passes through the point (1,-3) intersects the x-axis at a point. What are the coordinates of this point?

OpenStudy (raden):

the general equation of line is : y = mx + c with m=slope c=constant the first, subtitu value m and the (1,-3) to get c y = mx + c y = 5x + c -3 = 5(1) + c keep going...

OpenStudy (raden):

-3 = 5 + c find c?

OpenStudy (raden):

can u get it?

OpenStudy (anonymous):

Yeah I got it, thats why I closed the problem haha. Thanks.

OpenStudy (raden):

lol... yw:)